OpenStudy (anonymous):

If a hospital patient is given 50 milligrams of medicine which leaves the bloodstream at 10% per hour, how many milligrams of medicine will remain in the system after 6 hours? Use the function A(t) = Iert. 12.39 mg 27.44 mg 45.24 mg 91.11 mg

OpenStudy (anonymous):

HELP I REALLY NEED IT

OpenStudy (anonymous):

A(6) = 50e^(-.10*6) A(6) = 50e^(-.60) A(6) = 50e^(-.60) A(6) = 27.44 milligrams

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

good job yall are great